OpenStudy (anonymous):

Solve 5y^2/3+11y^1/3=-2

OpenStudy (lgbasallote):

could you put grouping symbols to avoid confusion??

OpenStudy (anonymous):

sorry new to this what do u mean?

OpenStudy (lgbasallote):

uhh like (5x^3)/(3+11) or something...because it can be interpreted in many ways

OpenStudy (anonymous):

ok sorry. (5y^2/3)+(11y^1/3)=-2

OpenStudy (lgbasallote):

ahh better...seems y^1/3 can be factored out...try taking it out...

OpenStudy (anonymous):

so i raise everything to the 3rd power? (125y^2)+1330y=-8 is that correct

OpenStudy (anonymous):

5y^2/3+11y^1/3=-2 let y^(1/3)=p 5p^2+11p+2=0